cynthia besch wants to buy a rug for a room that is 12 ft wide and 15 ft long. she wants to leave a uniform strip of floor around the rug. she can afford to buy 108 ft^2 of carpeting. what dimensions should the rug have?
Answer by mananth(16949)

cynthia besch wants to buy a rug for a room that is 12 ft wide and 15 ft long. she wants to leave a uniform strip of floor around the rug. she can afford to buy 108 ft^2 of carpeting. what dimensions should the rug have?
...
..
rug area = 108 ft^2
...
let width be x of the strip
..
width of rug = 12-2x
length of rug = 15-2x
Area of rug = (12-2x)(15-2x)=108
180-24x-30x+4x^2=108
4x^2-54x+72=0
/2
2x^2-27x+36=0
2x^2-24x-3x+36=0
2x(x-12)-3(x-12)=0
(x-12)(2x-3)=0
x= 12 OR = x=1.5 feet.
The width will be 1.5 feet.
m.ananth@hotmail.ca